OpenStudy (anonymous):

Find the most general antiderivative of the function. Use C for any needed constant. g(θ) = cos(θ) - 4sin(θ)

OpenStudy (anonymous):

y=int(g(x))=sin(x)+4cos(x) whic x=teta

OpenStudy (anonymous):

Don't forget to use capital 'C' to show there is a constant. Other than that, well done. :)

OpenStudy (anonymous):

oops sorry